Two numbers have the sum of 79. One number is 10 more than twice the other number. Find the values of the numbers?
Answer by addingup(3677)

Let's call the numbers x and 2x + 10. The problem says:
x + 2x + 10 = 79 Add x's on the left and subtract 10 on both sides
3x = 69
x = 23 and 2(23) + 10 = 56
Proof: 23 + 56 = 79